The sum of a number and three times another number is 18. Find the numbers if their product is a maximum. Please Help :)
Answer by scott8148(6628)

let x=number, y="other number" ___ x+3y=18 ___ x=18-3y

product=x*y ___ p=(18-3y)y ___ p=18y-3y^2

maximum is on the axis of symmetry ___ y=-b/2a ___ y=-18/(2(-3)) ___ y=3

x+3(3)=18 ___ x=9